2016 AMC 10A Problem 17

Attempt Problem 17 of the 2016 AMC 10A below, then check your answer against the professionally curated solution from LIVE by Po-Shen Loh. You can also try the full timed exam, view all 2016 AMC 10A solutions, or check the answer key.

All problems are used with official legal permission of the Mathematical Association of America (MAA).

17.

Let NN be a positive multiple of 5.5. One red ball and NN green balls are arranged in a line in random order. Let P(N)P(N) be the probability that at least 35\frac{3}{5} of the green balls are on the same side of the red ball. Observe that P(5)=1P(5)=1 and that P(N)P(N) approaches 45\frac{4}{5} as NN grows large. What is the sum of the digits of the least value of NN such that P(N)<321400?P(N) < \dfrac{321}{400}?
1212
1414
1616
1818
2020
Answer: A
Concepts:basic probabilityinequality
Small hint:
Place the red ball in one of the N+1N+1 gaps among the green balls
Big hint:
The bad gaps have between 2N5\frac{2N}{5} and 3N5\frac{3N}{5} green balls on the left
Solution:
Think of first arranging the NN green balls, then placing the red ball in one of the N+1N+1 gaps. If kk green balls are to the left of the red ball, then N−kN-k are to its right.
At least 35N\frac35N green balls are on one side exactly when k≤25Nk\le\frac25N or k≥35Nk\ge\frac35N. Thus the bad gaps are 25N+1,25N+2,…,35N−1,\frac25N+1,\frac25N+2,\ldots,\frac35N-1, a total of 15N−1\frac15N-1 gaps.
Therefore P(N)=1−N5−1N+1=4N+105N+5. \begin{aligned} P(N) &= 1-\frac{\frac{N}{5}-1}{N+1} \\ &= \frac{4N+10}{5N+5}. \end{aligned} Solving 4N+105N+5<321400\frac{4N+10}{5N+5}<\frac{321}{400} gives N>479N>479. The least positive multiple of 55 is 480480, whose digit sum is 1212.
Thus, the correct answer is A.
LIVE by Po-Shen Loh
Programs

LIVE Classes

Self-Paced Classes

Stars

Scholarship

About Us

Who Is Po-Shen Loh?

Email Contact

Newsletter

Jobs

Become a Star

For a more Thought Full world.

© 2026 Expii, Inc.